OpenStudy (anonymous):

What is the fourth root of 625x^10 ?

OpenStudy (anonymous):

\[\sqrt[4]{625x^ {10}}\]

OpenStudy (tkhunny):

5*5*5*5 = ??

OpenStudy (anonymous):

I got \[5x ^{2} \sqrt[4]{x^2}\]

OpenStudy (tkhunny):

You may wish to chop \(\sqrt[4]{x^{2}}\) down to \(\sqrt{|x|}\), but maybe not.